# Holographic entropy inequalities pass the majorization test

### v1

@@ Abstract
Quantities computed by minimal cuts, such as entanglement entropies achievable by the Ryu-Takayanagi proposal in the AdS/CFT correspondence, are constrained by linear inequalities. We prove a previously conjectured property of all such constraints: Any $k$ systems on the “greater-than” side of the inequality are subsumed in some $k$ systems on its “less-than” side (accounting for multiplicity). This finding adds evidence that the same inequalities also constrain the entropies under time-dependent conditions because it preempts a large class of potential counterexamples. We prove several other properties of holographic entropy inequalities and comment on their relation to quantum erasure correction and the Renormalization Group.
@@ 1 Introduction
The primary motivation for considering (2) comes from holographic duality or AdS/CFT correspondence adscft . In the presence of time reversal symmetry, von Neumann entropies of CFT regions are computed by equation (2), a statement known as the Ryu-Takayanagi (RT) proposal rt1 ; rt2 . In this context, the cuts are surfaces on the time reversal-symmetric spatial slice of the dual AdS spacetime. Equation (2) also applies to many tensor network states as an empirical observation meraref ; holoMERA ; HaPPYcode ; topoRTN , including random tensor networks rtn of arbitrary architecture where it holds as a theorem pagethm . Bounds, which are implied by prescription (2) are therefore necessary conditions for the emergence of a semiclassical bulk geometry as well as detectors of non-randomness of tensors in tensor network states. In common nomenclature, they are said to bound the holographic entropy cone hec .
@@ 2.1 A lightning review of contraction proofs
Matrix $X$ is an $N\times L$ matrix of 0’s and 1’s, which tell us whether $A_{p}$ ($1\leq p\leq N$) is contained in region $X_{i}$ ($1\leq i\leq L$ where $L$ counts LHS terms in the inequality.) Matrix $Y$ is defined analogously for the RHS; its dimensions are $N\times R$. Written in terms of matrix elements of $X$ and $Y$, inequality (3) takes the form:
@@
of matrix $X$ play a special role. For obvious reasons, we call $\vec{x}\,^{p}$ the LHS indicator vector of $A_{p}$. Similarly, the row vectors of matrix $Y$, denoted $\vec{y}\,^{p}$, are RHS indicator vectors of $A_{p}$. In what follows we denote components of vectors with subscripts, e.g. $(\vec{x}\,^{p})_{i}=x^{p}_{i}$.
@@ Acknowledgements.
We thank Guglielmo Grimaldi, Matthew Headrick, Veronika Hubeny and Pavel Shteyner for sharing a draft of their paper combinatorial and for synchronizing the arXiv submissions of our manuscripts. We thank Ricardo Espíndola, Xiaoliang Qi, Sirui Shuai and Dachen Zhang for discussions. BC thanks the organizers of the workshops “Discussions on Quantum Spacetime” held at LMSI (Mumbai, India) and “AI for Physics” held at the Aspen Center for Physics while XLW thanks the organizers of the 20th Asian Winter School on Strings, Particles and Cosmology held at IISER Bhopal (India), where part of this work was completed. This research was supported by an NSFC grant number 12042505. BC dedicates this paper to Stella Christie, Vice Minister of Higher Education, Science, and Technology of the Republic of Indonesia, on the occasion of her birthday.

### v2

@@ Abstract
Quantities computed by minimal cuts, such as entanglement entropies achievable by the Ryu-Takayanagi proposal in the AdS/CFT correspondence, are constrained by linear inequalities. We prove a previously conjectured property of all such constraints: Any $k$ systems on the “greater-than” side of the inequality whose overlap is nonempty are subsumed in some $k$ systems on its “less-than” side (accounting for multiplicity). This finding adds evidence that the same inequalities also constrain the entropies under time-dependent conditions because it preempts a large class of potential counterexamples. We prove several other properties of holographic entropy inequalities and comment on their relation to quantum erasure correction and the Renormalization Group.
@@ 1 Introduction
The primary motivation for considering (2) comes from holographic duality or AdS/CFT correspondence adscft . In the presence of time reversal symmetry, von Neumann entropies of CFT regions are computed by equation (2), a statement known as the Ryu-Takayanagi (RT) proposal rt1 ; rt2 . In this context, the cuts are surfaces on the time reversal-symmetric spatial slice of the dual AdS spacetime. Equation (2) also applies—sometimes up to a gauge subtlety—to many tensor network states as an empirical observation meraref ; holoMERA ; HaPPYcode ; xiwayne ; topoRTN ; vijaycharlie , including random tensor networks rtn of arbitrary architecture where it holds as a theorem pagethm . Bounds, which are implied by prescription (2) are therefore necessary conditions for the emergence of a semiclassical bulk geometry as well as detectors of non-randomness of tensors in tensor network states. In common nomenclature, they are said to bound the holographic entropy cone hec .
@@ 2.1 A lightning review of contraction proofs
Matrix $\mathcal{X}$ is an $N\times L$ matrix of 0’s and 1’s, which tell us whether $A_{p}$ ($1\leq p\leq N$) is contained in region $X_{i}$ ($1\leq i\leq L$ where $L$ counts LHS terms in the inequality.) Matrix $\mathcal{Y}$ is defined analogously for the RHS; its dimensions are $N\times R$. Written in terms of matrix elements of $\mathcal{X}$ and $\mathcal{Y}$, inequality (3) takes the form:
@@
of matrix $\mathcal{X}$ play a special role. For obvious reasons, we call $\vec{x}\,^{p}$ the LHS indicator vector of $A_{p}$. Similarly, the row vectors of matrix $\mathcal{Y}$, denoted $\vec{y}\,^{p}$, are RHS indicator vectors of $A_{p}$. In what follows we denote components of vectors with subscripts, e.g. $(\vec{x}\,^{p})_{i}=x^{p}_{i}$.
@@ Acknowledgements.
We thank Guglielmo Grimaldi, Matthew Headrick, Veronika Hubeny and Pavel Shteyner for sharing a draft of their paper combinatorial and for synchronizing the arXiv submissions of our manuscripts. We thank Charlie Cummings, Ricardo Espíndola, Xiaoliang Qi, Sirui Shuai and Dachen Zhang for discussions. BC thanks the organizers of the workshops “Discussions on Quantum Spacetime” held at LMSI (Mumbai, India) and “AI for Physics” held at the Aspen Center for Physics while XLW thanks the organizers of the 20th Asian Winter School on Strings, Particles and Cosmology held at IISER Bhopal (India), where part of this work was completed. This research was supported by an NSFC grant number 12042505. BC gives special thanks to Stella Christie, Vice Minister of Higher Education, Science, and Technology of the Republic of Indonesia, for the care and support given in the final stages of preparation of this manuscript.